# Dyonic RN-like and Taub-NUT-like black holes in Einstein-bumblebee gravity

### v1

@@ V.3 Unnormal thermodynamic quantities
Although the calculations are involved, the evaluation of the electric and magnetic potentials and charges follows standard procedures. In contrast, the computation of the mass and NUT charge remains debated. In this work, we employ the method of [105] to provide a set of Taub–NUT black hole thermodynamic quantities that satisfy the first law.
@@
$\Psi$ $=$ $-\frac{\sqrt{\ell+1}}{\kappa N(\ell+2)}\Bigg\{q-q\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r}\big)\Big]$ (121)
$+p\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r}\big)\Big]\Bigg\}+c_{m}\,$,
@@
The NUT parameter can give rise to corresponding electric and magnetic charges
$Q_{eN}$ $=$ $\int_{r=r_{h}}^{r\to\infty}\mathcal{F}_{tr}dr=2w_{2}\mathcal{A}_{t}|_{r\to\infty}^{r=r_{h}}$ (127)
$=$ $\frac{w_{2}\sqrt{\ell+1}}{\kappa N(\ell+2)}\Bigg\{q-q\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]$ (129)
$+p\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]\Bigg\}\,$,
$Q_{mN}$ $=$ $\int_{r=r_{h}}^{r\to\infty}F_{tr}dr=2w_{2}A_{t}|_{r\to\infty}^{r=r_{h}}$ (130)
$=$ $\frac{w_{2}}{N}\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\Bigg\{-p+p\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]$ (132)
$+q\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]\Bigg\}\,.$
@@
$=$ $\frac{1}{4}\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\Bigg\{-p+p\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]$ (135)
$+q\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]\Bigg\}\,$,
@@
$=$ $-\frac{\sqrt{\ell+1}}{4\kappa(\ell+2)}\Bigg\{q-q\cos\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]$ (138)
$+p\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]\Bigg\}\,.$
We find that there is a relationship between the electric and magnetic charges (113, 119) and the charges induced by the NUT parameter
$Q_{e}=\frac{w_{2}\sqrt{\ell+1}}{\kappa(\ell+2)}q-2\Phi_{N}Q_{eN}\,,\qquad Q_{m}=w_{2}\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}p+2\Phi_{N}Q_{mN}\,.$ (139)
@@
$-\delta\mathcal{H}_{\infty}=\frac{w_{2}\sqrt{\ell+1}}{\kappa}\delta\mu\,.$ (140)
However, as noted in [105], if the above expression were taken as the mass, the black hole mass could be either positive or negative. Since a non-positive-definite mass has no physical meaning, the expression represents only a partial contribution to the total mass. We find that the charges (132) and the potentials (138) induced by the NUT parameter can be consistently incorporated into the Wald formalism
$-\delta\mathcal{H}_{\infty}-\mathcal{H}_{T_{N}}-\delta\mathcal{H}_{T_{S}}=\delta M-\Phi_{N}\delta Q_{N}-\Phi_{eN}\delta Q_{eN}-\Phi_{mN}\delta Q_{mN}\,.$ (141)
@@
$M$ $=$ $\frac{w_{2}\sqrt{\ell+1}}{\kappa}\mu+2\Phi_{N}Q_{N}\,$, (142)
$Q_{N}$ $=$ $\frac{2w_{2}N\sqrt{\ell+1}}{\kappa r_{h}}\Bigg\{1-\frac{p^{2}+q^{2}}{4N^{3}(\ell+2)}\Big(2N$ (144)
$-r_{h}\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\sin\Big[2\sqrt{\frac{(\ell+1)(\ell+2)}{3\ell+2}}\tan^{-1}\big(\frac{N}{r_{h}}\big)\Big]\Big)\Bigg\}\,.$
@@ VII Conclusion
In this paper, we construct four-dimensional dyonic RN-like black holes with general topological horizons within Einstein-bumblebee gravity, one of the simplest vector-tensor theories realizing spontaneous Lorentz symmetry breaking. We then investigate their thermodynamic properties and employ the Wald formalism to compute the conserved mass and entropy, thereby establishing the first law of black hole thermodynamics. Furthermore, we extend the static dyonic solution to the Taub–NUT case, revealing its nontrivial black hole thermodynamics. Finally, we generalize the four-dimensional dyonic RN-like results to higher dimensions. These results provide a concrete framework for further studies of the effects of spontaneous Lorentz symmetry breaking in black hole physics and astrophysics.

### v2

@@ V.3 Unnormal thermodynamic quantities
Although the calculations are involved, the evaluation of the electric and magnetic potentials and charges follows standard procedures. In contrast, the computation of the mass and NUT charge remains debated. In this work, we employ the method of [105] to provide a set of Taub-NUT black hole thermodynamic quantities that satisfy the first law.
@@
$\Psi$ $=$ $-\frac{\sqrt{\ell+1}}{N(\ell+2)}\Big[q-q\cos X+p\sin X\Big]-c_{m}\,$, (107)
@@
The NUT parameter can give rise to corresponding electric and magnetic charges. Since the integrals associated with the NUT parameter correspond to the paths $T_{N,S}$ in Fig. 1, where $u=\pm 1$, the $dt\wedge dr$ components in (103) and (105) contribute to the two NUT-related electric and magnetic charges
$Q_{eN}$ $=$ $\frac{w_{2}}{\kappa}\int_{r_{h}}^{\infty}\mathcal{Q}_{tr}dr=\frac{w_{2}}{\kappa}\mathcal{A}_{t}|_{r\to\infty}^{r=r_{h}}$ (113)
$=$ $\frac{w_{2}\sqrt{\ell+1}}{\kappa N(\ell+2)}\Big[q-q\cos X(r_{h})+p\sin X(r_{h})\Big]\,$, (114)
$Q_{mN}$ $=$ $\frac{w_{2}}{\kappa}\int_{r_{h}}^{\infty}\mathcal{P}_{tr}dr=\frac{w_{2}}{\kappa}A_{t}|_{r\to\infty}^{r=r_{h}}$ (115)
$=$ $\frac{w_{2}}{2\kappa N}\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\Big[q\sin X(r_{h})+p\cos X(r_{h})-p\Big]\,.$ (116)
@@
$=$ $\frac{1}{4}\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\Big[q\sin X(r_{h})+p\cos X(r_{h})-p\Big]\,$, (118)
@@
$=$ $-\frac{\sqrt{\ell+1}}{2(\ell+2)}\Big[q-q\cos X(r_{h})+p\sin X(r_{h})\Big]\,.$ (120)
This suggests a relationship between the charges and potentials $Q_{eN}\Phi_{eN}+Q_{mN}\Phi_{mN}=0$ induced by the NUT parameter. We find that there is a relationship between the electric and magnetic charges (104, 106) and the charges induced by the NUT parameter
$Q_{e}$ $=$ $\frac{w_{2}\sqrt{\ell+1}}{\kappa(\ell+2)}q-2\Phi_{N}Q_{eN}\,$, (121)
$Q_{m}$ $=$ $\frac{w_{2}}{2\kappa}\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}p+2\Phi_{N}Q_{mN}\,.$ (122)
@@
$-\delta\mathcal{H}_{\infty}=\frac{w_{2}\sqrt{\ell+1}}{\kappa}\delta m\,.$ (123)
However, as noted in [105], if the above expression were taken as the mass, the black hole mass could be either positive or negative. Since a non-positive-definite mass has no physical meaning, the expression represents only a partial contribution to the total mass. This implies that the integrals over $T_{N,S}$ contribute partially to the total mass
$\delta\mathcal{H}_{T_{N}}+\delta\mathcal{H}_{T_{S}}=-\int_{r_{h}}^{\infty}H(r,u)\big|_{u=-1}^{u=1}dr\,$,
$H=\frac{\delta pu\sqrt{\ell+1}}{\kappa(\ell+2)(r^{2}+n^{2})^{2}}\Big[(r^{2}+n^{2})(p\cos X-q\sin X)+p(n^{2}-r^{2})\Big]$
$+\frac{\delta qu\sqrt{\ell+1}}{\kappa(\ell+2)(r^{2}+n^{2})^{2}}\Big[(r^{2}+n^{2})(q\cos X+p\sin X)+q(n^{2}-r^{2})\Big]-\frac{2n^{2}ru\sqrt{\ell+1}}{\kappa(r^{2}+n^{2})^{2}}\delta m$
$-\frac{\delta nu\sqrt{\ell+1}(p^{2}+q^{2})}{\kappa n(\ell+2)\sqrt{3\ell+2}(r^{2}+n^{2})^{2}}\Bigg[\sqrt{3\ell+2}(n^{2}-r^{2})\cos X+2nr\sqrt{(\ell+1)(\ell+2)}\sin X\Bigg]$
$-\frac{\delta nu\sqrt{\ell+1}}{\kappa n(\ell+2)(r^{2}+n^{2})^{3}}\Bigg\{r^{4}(p^{2}+q^{2})-3n^{2}r^{2}\Big[2r(\ell+2)(r-m)+p^{2}+q^{2}\Big]$
$+2n^{4}(\ell+2)\Big[r(6r-m)+n^{2}\Big]\Bigg\}\,.$ (124)
We find that the charges (116) and the potentials (120) induced by the NUT parameter can be consistently incorporated into the Wald formalism
$-\delta\mathcal{H}_{\infty}-\delta\mathcal{H}_{T_{N}}-\delta\mathcal{H}_{T_{S}}=\delta M-\Phi_{N}\delta Q_{N}-\Phi_{eN}\delta Q_{eN}-\Phi_{mN}\delta Q_{mN}\,.$ (125)
@@
$M$ $=$ $\frac{w_{2}\sqrt{\ell+1}}{\kappa}m+2\Phi_{N}Q_{N}\,$, (126)
$Q_{N}$ $=$ $\frac{2w_{2}N\sqrt{\ell+1}}{\kappa r_{h}}\Bigg\{1-\frac{p^{2}+q^{2}}{4N^{3}(\ell+2)}\Big[2N-r_{h}\sqrt{\frac{3\ell+2}{(\ell+1)(\ell+2)}}\sin X(r_{h})\Big]\Bigg\}\,.$ (127)
@@
Before concluding this section, we find it appropriate to clarify an important point. To date, a fully satisfactory and universally accepted understanding of the thermodynamics of Lorentzian Taub-NUT spacetimes has not yet been established. In Ref. [126], the authors employed the quasilocal action method and then computed all thermodynamic quantities using standard statistical mechanics techniques. They also recognized that the introduction of the NUT parameter gives rise to new charges in the thermodynamics, which [126] referred to as horizon electric and magnetic charges. These are precisely the $Q_{eN}$ and $Q_{mN}$ calculated in (116). Nevertheless, within that framework, it remains somewhat unclear whether these charges themselves, or instead their conjugate potentials, should be regarded as the appropriate independent thermodynamic variables, and a definitive resolution was not reached.
The literatures [127, 128] also computed the thermodynamics using the Wald formalism. However, [127] suggested that the mass should be proportional to the parameter $m$. Yet, the Taub–NUT solution exhibits a peculiar feature: it describes a black object with an event horizon for all real values of $m$. In other words, for nonzero $n$, a horizon exists for every $m\in(-\infty,\infty)$. In [128], the NUT potential is defined as being inversely proportional to the NUT parameter $N$, i.e., $\psi_{N}=1/(8\pi N)$, which implies that this thermodynamic framework lacks a well-defined $N\to 0$ limit. However, from the viewpoint of the metric, the Taub–NUT solution smoothly reduces to the Schwarzschild black hole as $N\to 0$. Hence, this suggests that further clarification of this thermodynamic interpretation may be desirable.
Similar discussions can be found in [129, 130, 131, 132, 133, 134, 135], where complete thermodynamic frameworks satisfying the first law are presented. The approach used in this work follows essentially the method proposed in [105]. Despite the variety of perspectives, one fact is clear: the introduction of the NUT parameter adds a new pair of thermodynamic conjugates to the black hole. If the black hole carries electric and magnetic charges, the NUT parameter also introduces the corresponding pair $Q_{e,mN}$ and $\Phi_{e,mN}$. Although the number of thermodynamic variables increases, the integration constants remain unchanged, implying that $Q_{e,mN}$ and $\Phi_{e,mN}$ are not independent. Moreover, for a rotating black hole, differences between the north and south poles emerge, and the thermodynamic quantities associated with the NUT parameter split accordingly.
Overall, these considerations indicate that the thermodynamics of Lorentzian Taub-NUT spacetimes remains a subtle and actively investigated subject. The successful application of our approach to the dyonic Taub-NUT black hole in Einstein-bumblebee gravity, viewed as a representative vector-tensor gravity theory, provides a concrete and consistent example within this broader context, and may serve as a useful reference for future studies of NUT-charged solutions in modified gravity theories.
@@ VII Conclusion
In this paper, we construct four-dimensional dyonic RN-like black holes with general topological horizons within Einstein-bumblebee gravity, one of the simplest vector-tensor theories realizing spontaneous Lorentz symmetry breaking. We then investigate their thermodynamic properties and employ the Wald formalism to compute the conserved mass and entropy, thereby establishing the first law of black hole thermodynamics. Furthermore, we extend the static dyonic solution to the Taub-NUT case, revealing its nontrivial black hole thermodynamics. Finally, we generalize the four-dimensional dyonic RN-like results to higher dimensions. These results provide a concrete framework for further studies of the effects of spontaneous Lorentz symmetry breaking in black hole physics and astrophysics.
Given that Einstein-bumblebee gravity admits GR-like exact black hole solutions involving an increasingly rich set of physical parameters, including mass, electric charge, magnetic charge, and the NUT charge, this highlights the advantages of employing this framework to investigate spontaneous Lorentz symmetry breaking. On the other hand, as emphasized in section III, although imposing a vanishing potential fixes the functional form of the bumblebee vector $B_{\mu}$ and reduces the resulting gravitational field equations to second order, the theory remains intrinsically a higher-derivative gravity. As in other higher-derivative gravity theories, the fact that one of the bulk geometries resembles—or even coincides with—its GR counterpart (for instance, many higher-curvature gravities admits the Schwarzschild black hole as an exact solution [136, 137, 138, 139, 140, 141]) does not imply that the boundary sector must exhibit the same trivial structure. Indeed, the computation of the Noether mass charge via the Wald formalism already demonstrates a clear distinction between Einstein-bumblebee gravity and GR at the boundary level. This motivates extending the present black hole solutions to include a cosmological constant and, through holographic methods, exploring the potential new effects such boundary differences may induce in the dual conformal field theory. Such an analysis could provide new insights into the structure and physical implications of vector-tensor theories of gravity.
Additionally, in a recent work [142], the authors investigated the implications of the Noether mass for observational constraints on spontaneous Lorentz violation, using another Lorentz-violating gravity—Einstein-Kalb-Ramond gravity—as an illustrative example. As shown in that analysis, although different mass definitions lead only to subleading corrections in weak-field observational tests of the Lorentz-violating parameter, the situation would be significantly different in strong-field regimes, such as those involving black holes or neutron stars where relativistic effects are pronounced. In such contexts, constraining the Lorentz-violating parameter using the Noether mass may yield non-negligible deviations compared with previous studies that relied predominantly on the Komar mass or the Arnowitt-Deser-Misner mass.